OpenStudy (anonymous):

Help me solve: A table for y = x^2 - 25 is given below. Solve each equation or inequality. x = -6, -5, -1, 0, 1, 5, 6 y = 11, 0, -24, -25, -24, 0, 11 a.) x^2 - 25 = 0 b.) x^2 - 25 < 0 c.) x^2 - 25 > 0

OpenStudy (e.mccormick):

Difference of squares, a stock equation, is what Luis popped in with, which is a good thing to remember!

OpenStudy (e.mccormick):

Take any two squares, like the \(x^2\) and \(b^2\) that he used, and their difference is always going to come out like that. \(x^2-b^2=(x-b)(x+b)\)

OpenStudy (anonymous):

Use difference of squares:\[(x^2-y^2)=(x-y)(x+y) \implies x^2-25 = (x-..)(x+..)\]Can you tell us what the expression turns in to now? @angelina22309

OpenStudy (anonymous):

Am I just suppose to plug in one of the x values it gives me?

OpenStudy (e.mccormick):

So what part of this one was giving you trouble?

OpenStudy (e.mccormick):

Ah, so they do not tell you what needs to be done in this one?

OpenStudy (anonymous):

Solve each equation or inequality

OpenStudy (e.mccormick):

AH. OK. Also, look at b and c, you put them in as the same.

OpenStudy (e.mccormick):

OK. So we have a), the \(x=\{-5,5\}\) and \(y=0\). At this point, I think this is asking you to account for al the numbers below, but not 100% sure. However, from finding that first thing you can work towards the others.

OpenStudy (anonymous):

I updated the question @e.mccormick

OpenStudy (e.mccormick):

OK. this looks like they want you to make it into a set of equations for when one thing or another is true.

OpenStudy (anonymous):

 Does this make sense? @angelina22309